OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line containing the longer diagonal of a quadrilateral whose vertices are A (2, 2), B(-2, -2), C(1, -1), and D(6, 4).

OpenStudy (anonymous):

Pleasee help me.

jimthompson5910 (jim_thompson5910):

how far did you get with this?

OpenStudy (anonymous):

I'm not sure where to begin.

jimthompson5910 (jim_thompson5910):

have you plotted the points?

jimthompson5910 (jim_thompson5910):

that's where I would start plot the points and connect them to form a quadrilateral

OpenStudy (anonymous):

Ok, after I plot them, what should I do after that.

jimthompson5910 (jim_thompson5910):

visually you should see which diagonal is longer are you able to see this?

OpenStudy (anonymous):

Yes.

jimthompson5910 (jim_thompson5910):

which diagonal is longer?

OpenStudy (anonymous):

CD

jimthompson5910 (jim_thompson5910):

a diagonal is not a side length (it passes through/near the center of the quadrilateral)

OpenStudy (anonymous):

Okay, I'm not sure which is longer than. I have it drawn on notebook paper, but I'm not sure.

jimthompson5910 (jim_thompson5910):

do you have geogebra? That's what I recommend to plot

OpenStudy (anonymous):

No I dont.

jimthompson5910 (jim_thompson5910):

alright, when you can, I recommend you download and install for now, I can just upload what I'm seeing

jimthompson5910 (jim_thompson5910):

see attached

jimthompson5910 (jim_thompson5910):

based on that, which diagonal is longer?

OpenStudy (anonymous):

The one that goes from point B to D?

jimthompson5910 (jim_thompson5910):

so segment BD

jimthompson5910 (jim_thompson5910):

sry i forgot the labels, but you are correct

jimthompson5910 (jim_thompson5910):

so you need to find the equation of the line that goes through B(-2, -2) and D(6, 4).

OpenStudy (anonymous):

Yes, how do I find the equation?

jimthompson5910 (jim_thompson5910):

do you know how to find the slope?

OpenStudy (anonymous):

No. I don't really understand the text in my assignments.

jimthompson5910 (jim_thompson5910):

what are the x and y coordinates of (-2, -2)

jimthompson5910 (jim_thompson5910):

same question about (6,4)

OpenStudy (anonymous):

I'm not sure what you mean by x and y coordinates. Wouldn't they be x (-2) .. y (-2) and then x (6) ... y (4) or are coordinates something else? I'm really confused. Sorry

jimthompson5910 (jim_thompson5910):

yeah for (-2,-2) we have x = -2 y = -2 also for (6,4), we have x = 6 y = 4

jimthompson5910 (jim_thompson5910):

to find the slope, we subtract the y coordinates 4 - (-2) = 4 + 2 = 6 then we subtract the x coordinates (same order) 6 - (-2) = 6 + 2 = 8

jimthompson5910 (jim_thompson5910):

After we get those differences, we divide the ydifference by the xdifference ydifference/xdifference = 6/8 = 3/4

jimthompson5910 (jim_thompson5910):

so the slope is 3/4

OpenStudy (anonymous):

How would put the equation in standard form then?

jimthompson5910 (jim_thompson5910):

have you seen something like y = mx+b before?

OpenStudy (anonymous):

Yes.

jimthompson5910 (jim_thompson5910):

what does m represent?

jimthompson5910 (jim_thompson5910):

it's ok to not know the answer, but please respond in some way

OpenStudy (anonymous):

I moved onto another question, sorry. I'll just ask my teacher tomorrow. Thanks tho! I appreciate your help! :)

jimthompson5910 (jim_thompson5910):

No worries and that's probably the best way to do it.